### v1

@@ Abstract
This paper presents a design for a drop tower test to evaluate a numerical model for a structurally reconfigurable spacecraft with actuatable joints, referred to as a transformable spacecraft. A mock-up robot for a 3U-sized transformable spacecraft is designed to fit in a limited time and space of the microgravity environment available in the drop tower. The robot performs agile reorientation, referred to as nonholonomic attitude control, by actuating joints in a particular manner. To adapt to the very short duration of microgravity in the drop tower test, a successive joint actuation maneuver is optimized to maximize the amount of attitude reorientation within the time constraint. The robot records the angular velocity history of all four bodies, and the data is analyzed to evaluate the accuracy of the numerical model. We confirm that the constructed numerical model sufficiently replicates the robot’s motion and show that the post-experiment model corrections further improve the accuracy of the numerical simulations. Finally, the difference between this drop tower test and the actual orbit demonstration is discussed to show the prospect.
@@
Drop tower test, Reconfigurable spacecraft, Attitude reorientation, Nonholonomic system, Particle swarm optimization, CubeSat1 Introduction
@@
Novel space exploration technologies have grown rapidly and widely in recent years. A transformable spacecraft is one such architecture, enabling a spacecraft to reconfigure itself with actuatable joints. We can expect the transformable spacecraft to perform multi-objective operations with highly redundant actuators. The Transformer research group, organized by JAXA and universities in Japan, has proposed and studied a concept of such a highly redundant ”Transformer” [15]. The other group, Tokyo Institute of Technology, launched HIBARI in 2021 [16, 5]. HIBARI had only four degrees of freedom actuators, but it carried out important and insightful orbit demonstrations with actuatable mechanisms.
@@
One key technology for the transformable spacecraft is ”nonholonomic attitude control,” where the spacecraft can reorient its attitude by successive joint actuation (the details are described in Section 3.1). Physical properties such as mass, position of the center of mass, and moment of inertia determine the direction and amount of the reorientation. Therefore, understanding the fidelity and accuracy of the numerical model is important for the maneuver design. In general, a spacecraft is composed of many components and is not necessarily precisely modeled due to time and cost constraints. Moreover, some components, such as flexible harnesses and propellants, take much work to model. Thus, measuring such physical properties through ground tests and operations in orbit is essential.
@@
This article describes a method of model evaluation using a drop tower facility. We introduce the experiment’s design and analyze the obtained results. Section 2 describes the facility of the drop tower, the design of a transformable robot, and a release mechanism. Section3 first introduces an overview of the nonholonomic attitude control and is followed by physical modeling and maneuver designs using the particle swarm optimization. Section 4 explains setups of the experiment and is followed by results and discussions in Section 5 and 6.
@@ 2.1 Facility of the drop tower
A drop tower is less costly than an orbit demonstration and a parabolic flight by an airplane. In addition, the test hardware can be refurbished on the ground during relatively long intervals of capsule drops. For these reasons, we adopted the drop tower as a test facility for this experiment. In particular, the drop tower facility COSMOTORRE in Hokkaido, Japan, was chosen for this experiment. The tower and the drop capsule are shown in Fig. 1, and their specification is overviewed in Table 1. Instruments for the experiment are equipped in the inner capsule. The inner capsule is installed inside the outer capsule, which reduces air drag during free fall.
@@ 2.2 Design of the transformable robot
The overview and dimensions of the designed robot are shown in Fig. 2 and 3. The robot is composed of four bodies with three actuatable joint units. Each joint unit comprises a servomotor, two absolute angle sensors (at both ends of the rotational axis of the motor), and bracket structures. The total mass is 2,847 g. The frame structure is composed of universal aluminum frames that can endure the landing impact at the end of the drop test. The robot’s size is compatible with 3U CubeSat, whereas the length in -direction is elongated due to the release mechanism.
@@
[…]
Two lasers are mounted orthogonally on body 2. The laser lay is projected on graph paper attached to the side walls of the test space and captured by cameras, which allows us to estimate the release velocity of the robot.
The robot has two microprocessors. Arduino micro mounted on the sequence control board controls the entire maneuver sequence of the robot. On the other hand, Teensy 4.1, mounted on the data handling board, specializes in data acquisition from MEMS gyroscopes and absolute angle sensors and data recording on an SD card. This separation of the roles enables the robot to achieve a sampling rate of about 280 Hz, which is sufficiently fast in this experiment.
@@ 2.3 Design of the release mechanism
To push the robot with the proper vertical velocity at releasing
@@
The robot is attached to the electromagnet unit with the magnetic plate of SUS440C. The electromagnet, Kanetec KE-3HA, provides a magnetic force of 200 N (20.4 kgf) at a maximum, which can sufficiently hold the robot weighing about 3 kg. However, an extra holding mechanism with nylon wires shown in Fig. 5 is attached because higher acceleration is added when a crane lifts the capsule. After crane lifting is completed, the nylon wire is tied firmly between hooks and burned by a nichrome electric heating wire. Tensile force on the nylon wires is applied by sliding the position of wire tensioners, which enables tight holding and secure cutting. Requirement 1, shown above, is satisfied with this electromagnet and nylon wire holding mechanism.
The magnetic plate is connected to the shaft with a spring and fits in the 3D-printed housing structure. The spring is compressed when the robot is attached, and thus, it provides detaching force to overcome residual magnetic force under micro-gravity. The residual magnetic force is estimated as 5% of the maximum magnetic force, hence 10 N, according to the datasheet. The designed spring mechanism provides 11.1 N, which can give sufficient detaching force under microgravity. In addition, the stroke of the spring is set to 4 mm, which can give sufficient separation distance (1 mm is enough according to the datasheet). Requirement 2 is satisfied with this mechanism.
Two release detection switches (OMRON SS-5GL) are equipped beside the magnetic plate. The switches are closed when the robot is attached to the electromagnet and opened when released. The robot starts a preset sequence after detecting the switching. In addition, the switches slightly push off the robots at the release and provide proper vertical velocity. If the vertical velocity is small, the robot must wait very long until the beginning of movement to prevent itself from colliding with the magnetic releaser. On the other hand, if the vertical velocity is too large, the robot collides with the bottom of the test space before the end of the attitude maneuver. Thus, the proper vertical velocity range is estimated as mm/s.
@@ 3 Maneuver design
This section describes the design of the reorientation maneuver adopted in the experiment. First, we overview the principles of attitude reorientation and a fundamental maneuver in Section 3.1. Section 3.2 denotes the physical modeling of the robot.
@@ 3.1 Nonholonomic attitude reorientation
The dynamics of free-floating multibody have been intensively investigated. It is known that the free-flying multibody can reorient itself by a proper sequence of joint actuation, even if the total angular momentum has always been kept at zero. Historically, it is known by righting reflex of falling cats [9, 7, 3], and intensively studied for space robots with robotic manipulators [10, 1, 11, 13, 8]. The angular momentum conservation law imposed on the free-flying robot is nonholonomic, which means the robot’s final attitude depends on the procedure of joint actuation. Therefore, the robot can arbitrarily reorient its attitude by designing proper joint actuation maneuvers.
In this experiment, we adopt a maneuver similar to the one proposed in previous research, such as [12, 6]. Two examples of joint actuation are shown in Fig. 8. The sequence shown in the upper side is composed of four consecutive joint actuations, where the first stroke and third strokes are actuations of the same joint but in the opposite direction ( and ), and so do the second and fourth strokes ( and ). The attitude is reoriented due to the nonholonomy of the system, and thus it is referred to as nonholonomic attitude reorientation. The one in the lower is composed of four consecutive joint actuation, but the first and fourth strokes are the same pattern as the second and third strokes. This sequence exactly cancels the intermediate attitude difference, and therefore, attitude is not reoriented after the four strokes. These two examples show that the final attitude depends on the intermediate joint actuation sequence.
The corresponding trajectory in a joint angle space is shown in the left side of Fig. 8. A joint angle space is a multi-dimensional space where each point corresponds to one set of joint states. In these examples, the maneuvers are expressed as a consecutive rectilinear path. If the path draws a closed polygon as in the upper case in Fig. 8, the joint state of a robot returns to the initial state, whereas attitude is changed from the initial. This closed-loop actuation is adopted in this experiment because it clearly shows the effect of reorientation due to the nonholonomy of the system.
@@ 3.2 Physical modeling of the robot
This section describes how to model physical properties such as center of mass, moment of inertia, and joint positions of joints. These properties are necessary to design optimal reorientation maneuvers in Section 3.3.
@@
Figure 9 shows the definitions of global body frames and local body frames, indicating the origins and orientations of the frames. The origins are attached to the corner point of the aluminum frame. Table 4 shows the positions of the rotational axes of joints in the global frame. Note that the axes are aligned parallel to or in the global frame. Therefore, for example, the position of joint 1 is not indicated since the axis is parallel to the global axis. Table 5 shows the positions of the center of mass of each body in the local body frame. Table 6 is the moment of inertia of each body around the local center of mass in the local body frame.
Joint 1 | - | 9.9 | 236.0 |
Joint 2 | 90.1 | - | 112.0 |
Joint 3 | - | 112.0 | 10.0 |
[…]
@@ 3.3 Design of optimal reorientation maneuver
In this experiment, we adopted the particle swarm optimization (PSO). PSO is a bio-inspired heuristic optimization method proposed in [4]. The formulation is simple and requires no gradient information for the objective function. Moreover, it is easily extensible to the robot with more actuatable joints and, thus, directly applicable in future missions.
The joint actuation sequence is expressed as a closed loop in a joint angle space as described in Section 3.1. A schematic of an example of joint actuation sequence is shown in Fig. 10. A point in the joint angle space corresponds to a set of specific joint angles. Angles of joints 1, 2, and 3 are denoted as whereas input joint angle displacements are denoted as respectively. Specifically, Fig. 10 shows the case of , . For a certain set of , the total number of combinations of joint actuation sequence is 30, as shown in Fig. 11. Note that is not chosen just after because it does not induce any attitude difference.
[…]
@@ 3.4 Design of experiment sequence
After the pausing, the robot starts the maneuver designed in Section 3.3. The interval between the actuation is calculated as the joints move at 170 deg/s on average. After the four strokes, the robot pauses the joint actuation and continues MEMS gyroscope and joint angle data recording until the deceleration of the capsule.
@@
0 | 0.00 | Capsule dropped, electromagnet released, and release detected |
1 | 0.70 | Joint 1 commanded to be actuated ( degrees) |
2 | 1.05 | Joint 2 commanded to be actuated ( degrees) |
3 | 1.47 | Joint 1 commanded to be actuated ( degrees) |
4 | 1.82 | Joint 2 commanded to be actuated ( degrees) |
5 | 2.23 | Maneuver finished, and data recording maintained |
6 | 2.50 | Drop capsule decelerated at landing |
[…]

### v2

@@ Abstract
This paper presents a design for a drop tower test to evaluate a numerical model for a structurally reconfigurable spacecraft with actuatable joints, referred to as a transformable spacecraft. A mock-up robot for a 3U-sized transformable spacecraft is designed to fit within the limited time and space constraints of the microgravity environment available in the drop tower. The robot performs agile reorientation, referred to as nonholonomic attitude control, by actuating joints in a particular manner. To adapt to the very short duration of microgravity in the drop tower test, a successive joint actuation maneuver is optimized to maximize the amount of attitude reorientation within the time constraint. The robot records the angular velocity history of all four bodies, and the data is analyzed to evaluate the accuracy of the numerical model. We confirm that the constructed numerical model sufficiently replicates the robot’s motion and show that the post-experiment model corrections further improve the accuracy of the numerical simulations. Finally, the difference between this drop tower test and the actual orbit demonstration is discussed to show the prospect.
@@
[…]
@@
Novel space exploration technologies have grown rapidly and widely in recent years. Future space robots (including spacecraft in a broad sense) will need to be, for example, more autonomous, more intelligent, more adaptive, more dexterous, and more economical. A transformable spacecraft is one promising architecture, enabling a spacecraft to reconfigure itself with actuatable joints. We can expect the transformable spacecraft to perform multi-objective operations with highly redundant actuators. The Transformer research group, organized by JAXA and universities in Japan, has proposed and studied a concept of such a highly redundant "Transformer" [1]. The other group, Tokyo Institute of Technology, launched HIBARI in 2021 [2, 3]. HIBARI had only four degrees of freedom actuators, but it carried out important and insightful orbit demonstrations with actuatable mechanisms.
One key technology for the transformable spacecraft is "nonholonomic attitude control," where the spacecraft can reorient its attitude by successive joint actuation. The study of nonholonomic systems has a long history. Comprehensive surveys can be found in textbooks such as that by Bloch [4]. As for the mechanics of free-flyers, one of the most widely known nonholonomic problems is the righting reflex of falling cats (or the falling-cat phenomenon). This is a motion in which a cat can always land in an upright attitude by twisting its body in the air, even when dropped in an upside-down attitude. There have been many studies to elucidate the dynamics of this phenomenon [5][6]. Since the 1980s, when numerous space robots were proposed to construct space stations [7][8], more studies have focused on the dynamics of free-flying robots. Some researchers have proposed actively using the attitude reaction when a manipulator is actuated. Many attitude reorientation methods have been proposed that leverage the nonholonomy of the attitude motion [9][10][11][12][13][14][15][16][17].
@@
Physical properties such as mass, position of the center of mass, and moment of inertia determine the direction and amount of the reorientation of the nonholonomic attitude control. Therefore, understanding the fidelity and accuracy of the numerical model is important for the maneuver design. In general, a spacecraft is composed of many components and is not necessarily precisely modeled due to time and cost constraints. Moreover, some components, such as flexible harnesses and liquid propellants, cannot be modeled precisely. Thus, measuring such physical properties through ground tests and operations in orbit is essential.
Hardware-based experiments are essential to evaluate the modeling accuracy of the physical properties of robots. There are several examples of experiments to demonstrate the nonholonomic attitude control in hardware. Some researchers attempted to apply the righting reflex of falling cats to free-flying robots or spacecrafts [18][19][20]. The most active projects in recent years are carried out by ETH Zurich, named "SpaceBok" [21] and "SpaceHopper" [22]. Both are legged robots designed to walk and jump in a low-gravity space environment. After a jump on the surface, the robot changes its attitude by nonholonomic attitude control using its legs. To construct the reorientation control law, they adopted deep reinforcement learning in a simulation environment. The simulation parameters, such as masses, moment of inertia, and friction coefficients, cannot be precisely modeled in the simulator, and thus, they used the method of domain randomization [23]. The policy was trained on many environments in parallel, each having a randomized set of simulation parameters with 20–50% uncertainty. However, such uncertainty may deteriorate the capability of attitude control. In this situation, evaluating and updating such physical parameters using the obtained experimental data can contribute to enhancing attitude control capability.
@@
This article describes a method of model evaluation using a drop tower facility. We introduce the experiment’s design and analyze the obtained results. The motivation of the experiment is to estimate the modeling error and the performance of attitude reorientation. Moreover, by extrapolating the result obtained in the drop tower test, we can estimate the on-orbit reorientation performance assuming a certain 3U transformable CubeSat. Section 2 describes the facility of the drop tower, the design of a transformable robot, and a release mechanism. Section 3 first introduces an overview of the nonholonomic attitude control and is followed by physical modeling and maneuver designs using the particle swarm optimization. Section 4 explains setups of the experiment and is followed by results in Section 5. Finally, Section 6 discusses expected on-orbit performance and several differences between the ground and on-orbit experiment. The preprint version of this article can be found in the arXiv server [24].
@@ 2.1 Facility of the drop tower
A drop tower is less costly than an orbit demonstration and a parabolic flight by an airplane. In addition, the test hardware can be refurbished on the ground during relatively long intervals between capsule drops. For these reasons, we adopted the drop tower as a test facility for this experiment. In particular, the drop tower facility COSMOTORRE in Hokkaido, Japan, was chosen for this experiment. The tower and the drop capsule are shown in Fig. 1, and their specification is overviewed in Table 1. Instruments for the experiment are equipped in the inner capsule. The inner capsule is installed inside the outer capsule, which reduces air drag during free fall.
@@ 2.2 Design of the transformable robot
The overview and dimensions of the designed robot are shown in Fig. 2 and 3. The robot is composed of four bodies with three actuatable joint units. Each joint unit comprises a servomotor, two absolute angle sensors (at both ends of the rotational axis of the motor), and bracket structures. The total mass is 2,847 g. The frame structure is composed of universal aluminum frames that can endure the landing impact at the end of the drop test. The robot’s size is compatible with a 3U CubeSat, whereas the length in the -direction is elongated due to the release mechanism.
@@
Two lasers are mounted orthogonally on body 2. The laser ray is projected on graph paper attached to the side walls of the test space and captured by cameras, which allows us to estimate the release velocity of the robot.
The robot has two microprocessors. The Arduino micro, mounted on the sequence control board, controls the entire maneuver sequence of the robot. On the other hand, Teensy 4.1, mounted on the data handling board, specializes in data acquisition from MEMS gyroscopes and absolute angle sensors and data recording on an SD card. This separation of the roles enables the robot to achieve a sampling rate of about 280 Hz, which is sufficiently fast in this experiment.
@@ 2.3 Design of the release mechanism
To push the robot with the proper vertical velocity at release
@@
The robot is attached to the electromagnet unit with the magnetic plate of SUS440C. The electromagnet, Kanetec KE-3HA, provides a magnetic force of 200 N (20.4 kgf) at a maximum, which can sufficiently hold the robot weighing about 3 kg. However, an extra holding mechanism with nylon wires shown in Fig. 5 is attached because higher acceleration is added when a crane lifts the capsule. After the crane lifting is completed, the nylon wire is tied firmly between hooks and burned by a nichrome electric heating wire. Tensile force on the nylon wires is applied by sliding the position of wire tensioners, which enables tight holding and secure cutting. Requirement 1, shown above, is satisfied with this electromagnet and nylon wire holding mechanism.
The magnetic plate is connected to the shaft with a spring and fits in the 3D-printed housing structure. The spring is compressed when the robot is attached, and thus, it provides a detaching force to overcome residual magnetic force under micro-gravity. The residual magnetic force is estimated as 5% of the maximum magnetic force, hence 10 N, according to the datasheet. The designed spring mechanism provides 11.1 N, which can give sufficient detaching force under microgravity. In addition, the stroke of the spring is set to 4 mm, which can give sufficient separation distance (1 mm is enough according to the datasheet). Requirement 2 is satisfied with this mechanism.
Two release detection switches (OMRON SS-5GL) are equipped beside the magnetic plate. The switches are closed when the robot is attached to the electromagnet and opened when released. The robot starts a preset sequence after detecting the switch. In addition, the switches slightly push off the robots at the release and provide proper vertical velocity. If the vertical velocity is small, the robot must wait for a very long time until the beginning of movement to prevent itself from colliding with the magnetic releaser. On the other hand, if the vertical velocity is too large, the robot collides with the bottom of the test space before the end of the attitude maneuver. Thus, the proper vertical velocity range is estimated as mm/s.
@@ 3 Maneuver design
This section describes the design of the reorientation maneuver adopted in the experiment. First, we provide an overview of the principles of attitude reorientation and a fundamental maneuver in Section 3.1. Section 3.2 denotes the physical modeling of the robot.
@@ 3.1 Nonholonomic attitude reorientation
The dynamics of free-floating multibodies have been intensively investigated. It is known that the free-flying multibody can reorient itself by a proper sequence of joint actuation, even if the total angular momentum has always been kept at zero. Historically, it is known by the righting reflex of falling cats [5, 25, 6], and intensively studied for space robots with robotic manipulators [26, 27, 9, 28, 10]. The angular momentum conservation law imposed on the free-flying robot is nonholonomic, which means the robot’s final attitude depends on the procedure of joint actuation. Therefore, the robot can arbitrarily reorient its attitude by designing proper joint actuation maneuvers. In this experiment, we adopt a maneuver similar to the one proposed in previous research, such as [29, 30]. Two examples of joint actuation are shown in Fig. 8. The sequence shown on the upper side is composed of four consecutive joint actuations, where the first stroke and third strokes are actuations of the same joint but in the opposite direction ( and ), and so do the second and fourth strokes ( and ). The attitude is reoriented due to the nonholonomy of the system, and thus it is referred to as nonholonomic attitude reorientation. The one in the lower is composed of four consecutive joint actuations, but the first and fourth strokes are the same pattern as the second and third strokes. This sequence exactly cancels the intermediate attitude difference, and therefore, attitude is not reoriented after the four strokes. These two examples show that the final attitude depends on the intermediate joint actuation sequence.
The corresponding trajectory in a joint angle space is shown on the left side of Fig. 8. A joint angle space is a multi-dimensional space where each point corresponds to one set of joint states. In these examples, the maneuvers are expressed as a consecutive rectilinear path. If the path draws a closed polygon as in the upper case in Fig. 8, the joint state of a robot returns to the initial state, whereas the attitude is changed from the initial. This closed-loop actuation is adopted in this experiment because it clearly shows the effect of reorientation due to the nonholonomy of the system.
@@ 3.2 Physical modeling of the robot
This section describes how to model physical properties of the robot. The robot is modeled as a multi-rigid-body system, and thus, the required physical properties to describe the robot are the mass, the position of center of mass, the moment of inertia, directions of joints, and the positions of joints relative to the center of mass of adjacent bodies. The numerical model is necessary to design optimal reorientation maneuvers in Section 3.3.
@@
Figure 9 shows the definitions of global body frames and local body frames, indicating the origins and orientations of the frames. The origins are attached to the corner point of the aluminum frame. Table 4 shows the positions of the rotational axes of joints in the global frame and the local frame. The positions in the global frame are the values when all the joint angles are set to be zero. In addition, the local position of the -th joint is expressed in the local frame of the -th body. Note that the axes are aligned parallel to or in the global frame. Therefore, for example, the position of joint 1 is not indicated since the axis is parallel to the global axis. Table 5 shows the positions of the center of mass of each body in the local body frame. Table 6 is the moment of inertia of each body around the local center of mass in the local body frame. All these physical properties include modeling errors due to the assumptions described above. In particular, the center of mass position and the moment of inertia have relatively large modeling errors because they cannot be directly measured from the actual hardware. In this article, we attempt to compensate for the modeling error using the experimental data in Section 5.
| global | global | global | local | local | local | |
| Joint 1 | - | 10.0 | 236.0 | - | 10.0 | 112.0 |
| Joint 2 | 90.0 | - | 112.0 | 90.0 | - | 112.0 |
| Joint 3 | - | 112.0 | 10.0 | - | 5.0 | -25.0 |
@@ 3.3 Design of optimal reorientation maneuver
First, we set a constraint that only one joint can be moved in one stroke because the actuations of motors used in this experiment cannot be synchronized due to the hardware limitation. To identify a joint actuation sequence that induces a large attitude difference, we optimized joint angle displacements and the joint movement order separately. When optimizing the joint angle displacements, gradient-based optimization is difficult to apply because the objective function may be non-smooth in the design space. Therefore, we adopted the particle swarm optimization (PSO) to identify the joint angle displacements. PSO is a bio-inspired metaheuristic optimization method proposed in [31]. The formulation is simple and requires no gradient information for the objective function. Moreover, it is easily extensible to the robot with more actuatable joints and, thus, directly applicable in future missions.
In PSO, particles are prepared in the design space, and their initial positions and velocities are set randomly in the space. Then, the particles are updated according to the following rule [32]:
where and are the position and the velocity of the -th particle in the design space at the -th iteration, is the inertia weight, and are the learning factors. Large imposes relatively high importance on the experience of each particle, whereas large imposes relatively high importance on the experience of all particles. and are random numbers sampled uniformly from 0.0 to 1.0, and and are so far the best positions of the -th particle and all the particles at the -th iteration. Here, the position with the smallest objective function is selected as the best one. The updates are repeated until the number of iterations reaches . To achieve efficient exploration of particles in PSO, coefficients , , are linearly changed according to the number of iterations in the following rule [32]:
Note that we determined the constants in these equations through trial and error. In the above rules, the experience of each particle is given importance at the beginning of the update, and experiences of all particles are given more importance as the number of iterations increases. In addition, to balance the computational cost and the convergence speed of PSO, we set the number of particles to , and the number of iterations to .
The joint actuation sequence is expressed as a closed loop in a joint angle space as described in Section 3.1. A schematic of an example of joint actuation sequence is shown in Fig. 10. A point in the joint angle space corresponds to a set of specific joint angles. Angles of joints 1, 2, and 3 are denoted as , whereas input joint angle displacements are denoted as , respectively. Specifically, Fig. 10 shows the case of , . For a certain set of , the total number of combinations of joint actuation sequence is 30, as shown in Fig. 11. Note that is not chosen just after because it does not induce any attitude difference. In this experiment, we define the joint angle displacements as the design parameters . Then, we alternately perform optimizing the joint angles via PSO and the joint movement sequence for a given set of them heuristically.
The optimization problem is defined as Eq. (4). The objective function is evaluated by the magnitude of attitude difference and the penalty on maneuver time overrun . The penalty increases when , whereas it remains zero when . If is large, the maneuver time overrun is avoided more strictly. In this experiment, is chosen as a sufficiently large value. is set to 1.5 sec, assuming a 1.0 sec margin should be included in the total 2.5 sec duration.
Applying PSO to the free-floating robot defined in Section 3.2, we obtained the optimized maneuver as shown in Fig. 12. Optimized joint angle displacements are degrees. The average angular velocity of the joint is set to be 170 deg/s, resulting in 1.50 sec of the entire maneuver time. The resulting maneuver is composed only of 4 strokes because is degenerated into 0 degrees through the optimization. This is because the moment of inertia of body 4 is smaller than the others by one order, as in Table 6, and thus, the actuation of joint 3 is omitted in the limited maneuver time. The total attitude difference is 26.11 degrees in the axis. The resulting maneuver is tested in the drop tower test as shown in Section 5. In the actual test, the values of the joint angle displacements commanded to motors are rounded as to simplify the operations of the microprocessor.
@@ 3.4 Design of experiment sequence
After the pause, the robot starts the maneuver designed in Section 3.3. The interval between the actuations is calculated as the joints move at 170 deg/s on average. After the four strokes, the robot pauses the joint actuation and continues MEMS gyroscope and joint angle data recording until the capsule decelerates.
@@
| 0 | 0.00 | Capsule dropped, electromagnet released, and release detected |
| 1 | 0.70 | Joint 1 commanded to be actuated ( degrees) |
| 2 | 1.05 | Joint 2 commanded to be actuated ( degrees) |
| 3 | 1.47 | Joint 1 commanded to be actuated ( degrees) |
| 4 | 1.82 | Joint 2 commanded to be actuated ( degrees) |
| 5 | 2.23 | Maneuver finished, and data recording maintained |
| 6 | 2.50 | Drop capsule decelerated at landing |